Both power series solutions of

y'' + ln(x + 1)y' + y = 0

centered at the ordinary point

x = 0

are guaranteed to converge for all x in which one of the following intervals?

(−1, ∞)

1
2

,

1
2

     (−∞, ∞) [−1, 1]
Answer #1

į an (n) uhy Sely Given differential ean y + In pell) y + y = 0 Power Series & (2) hro y = { anth Now differentiate egn (2)Ant 2 (2+2) (n + 1) + In (2+1) anti (ht) + an/Hh In ]wetre T neo anez (h+2)(n+1) + ln (4+1) anti (n+1) + an ao - Ant 2 (h+2)Put na 3 as 93 (5) (4) - - az A5 11 = as 190 1 As 51 bring power Series 401) 8 andh 17 Qoto + Q, H & Agt? + Q 43 +99414 + aguY(1) = ao H 21 ? + *** + 7-abu ta, H-43 3 Ao Cubot a Sincen 7 4 () = ao Car (o) tai Sin (o) ( Hao given) Y. ao (1) + a, (o) 9